# Threshold-less and Flexibly Tunable Frequency Comb via Floquet Engineering

### v2

@@ Acknowledgments
We would like to express our gratitude for the fruitful discussions with Kaiye Shi, and Qiongyi He. Y. Liu acknowledges the support of Beijing Natural Science Foundation (Z240007), National Natural Science Foundation of China (Nos. 92365210 and 12374325), CAST Young Elite Scientists Sponsorship Program (Grant No. 2023QNRC001), and Beijing Municipal Science and Technology Commission (Grant No. Z221100002722011). This work was also supported by the Innovation Program for Quantum Science and Technology (2021ZD0302200), the Chinese Academy of Sciences (No. GJJSTD20200001), the National Key R&D Program of China (Grants No. 2021YFB3202800).
@@
S. W. conducted the measurements, Y. L. conceived the project and coordinated the research. All authors contributed to the analysis of the measurement results, wrote the manuscript, and provided their final approval for publication.

### v3

@@ Acknowledgments
The authors sincerely thank fruitful discussions with Kaiye Shi and Qiongyi He on the Floquet Hamiltonian.
Funding
Y.L. discloses support for the research of this work from Beijing Natural Science Foundation (No. Z240007), National Natural Science Foundation of China (No. 12374325 and No. 92365210), CAST Young Elite Scientists Sponsorship Program (No. 2023QNRC001), and Beijing Municipal Science and Technology Commission (No. Z221100002722011). J.Z. discloses support for publication of this work from the Innovation Program for Quantum Science and Technology (No. 2021ZD0302200), the Chinese Academy of Sciences (No. GJJSTD20200001), the National Key R&D Program of China (No. 2021YFB3202800). M.A.S. discloses support for publication of this work from European Research Council (contract 101019712), Research Council of Finland (Finnish Quantum Flagship, No. 358877). S.W., C.W., M.H.J.d.J., and L.M.d.L. declare no relevant funding.
@@
S.W. performed the experiments and carried out the theoretical analysis. C.W. fabricated the devices. M.H.J.d.J. and S.W. developed the numerical simulations. Y.L. conceived the idea, coordinated the project, and supervised the research. S.W., C.W., M.H.J.d.J., L.M.L., J.Z., M.A.S. and Y.L. contributed to the analysis of the measurement results, wrote the manuscript, and provided their final approval for publication.